60 ml of standard hardness containing 1 mg of pure CaCO3 per ml consumed 22 ml of EDTA. 40 ml 

of water sample consumed 20 ml of EDTA solution using EBT indicator. 40 ml of water sample after 

boiling, filtering consumed 15 ml of EDTA. Calculate the temporary and permanent hardness of 

water sample.


1
Expert's answer
2021-06-15T05:13:48-0400

Given:-

A numerical problem based on EDTA complexometric titration, which is to be accomplished.

"Hardness =\frac{Volume of EDTA consumed }{Volume of hardwater}×1000"

Temporary hardness   = Total hardness- Permanent hardness


To find the hardness of standard hardwater

   V1 mL of EDTA = V1 mg of CaCO3

Volume of EDTA consumed    = 22 mL

Volume of standard hard water = 60mL

"Hardness =\frac{22mL}{60mL}×1000"


=366.67 ppm


To find the hardness before boiling the sample water

Volume of EDTA consumed    = 20 mL

Volume of standard hard water = 40mL

"Hardness =\frac{20mL}{40mL}×1000"


= 500 ppm

Total hardness = 500 ppm


To find the hardness after boiling the sample water

After boiling the sample water temporary hardness is eliminated. Now, only the permanent hardness will be found.

Volume of EDTA consumed    = 15 mL

Volume of standard hard water = 40mL

"Hardness =\frac{15mL}{40mL}×1000"


Permanent hardness= 375 ppm

To find temporary hardness therefore,

Temporary hardness = Total hardness - Permanent hardnes

   = 500 - 375

      = 125 ppm

Permanent hardness = 375 ppm

Temporary hardness = 125 ppm